A propane cylinder having the volume of 20 liters, has been used outside in a temperature of 8 degrees c
And found to be virtually empty . The valve is closed and is left out in the sun where it's temperature increases to 40 degrees c determine;
A) the pressure of the cylinder at this temperature
B) the mass of propane still in the cylindery
1
Expert's answer
2012-06-19T09:16:00-0400
pV=(m/M)RT
T=8+273=281
T1=281+40=321
M=44
m=44
V=20
R=8.314
a)
p=mRT1/(VM)=8.314*321/20=133.44
b)
m=44 --- molecular mass